OpenStudy (anonymous):

Simplify the expression. Write your answer in scientific notation. (4 x 10^8)(8.2 x 10^-2)

OpenStudy (anonymous):

3.28×10^7

jimthompson5910 (jim_thompson5910):

Multiply 4 and 8.2 to get 4*8.2 = 32.8 Then multiply 10^8 and 10^(-2) by adding the exponents to get 10^8*10^(-2) = 10^(8 + (-2)) 10^8*10^(-2) = 10^(8 - 2) 10^8*10^(-2) = 10^6 --------------------------------- So we now have 32.8 x 10^6 Now convert 32.8 to scientific notation to get 3.28 x 10^1 --------------------------------- We now have (3.28 x 10^1) x 10^6 3.28 x (10^1 x 10^6) 3.28 x 10^(1 + 6) 3.28 x 10^7 So that means (4 x 10^8)(8.2 x 10^-2) = 3.28 x 10^7

OpenStudy (anonymous):

Thank you!! :)

jimthompson5910 (jim_thompson5910):

you're welcome